OpenStudy (anonymous):

factor: 128x^5+54x^2y^3

OpenStudy (zehanz):

This one is easier than the last one, becuase there are only two terms. Can you see numbers that are factors of both terms, 128x^5 and 54x^2y^3?

OpenStudy (anonymous):

ok and no.... im really bad at these....... :/

OpenStudy (zehanz):

128 = 2*64, 54 = 2*27 So 2 is a factor of 128 and of 54. This means you could write it as\[2(64x^5+27x^2y^3)\] But I can see more common factors now...(hint: x)

OpenStudy (anonymous):

i figured it out thank you!!

OpenStudy (zehanz):

\[2x^2(64x^3+27y^3)\]

OpenStudy (anonymous):

thats what i got thank you tons!! :)

OpenStudy (zehanz):

YW ;)